# ConceptDrift: Uncovering Biases through the Lens of Foundation Models

### v2

@@ Appendix C Algorithm
Output: Identified positively correlated biases $\mathcal{B}$.
@@
16:      $\mathcal{B}=\mathcal{B}\cup(k,b_{k})$ // Positively correlated biases
17:   end for
18:   return  $\mathcal{B}$
Appendix D Loss to bias similarity correlation
To validate that the environments discovered through the partitioning approach from B2T [15], using the biases uncovered with our method, are indeed effective in reducing the model’s reliance on shortcuts, we analyzed the correlation between the similarity of each training image to biases and its training loss value under both ERM and GroupDRO training, using our identified groups. The results show a reduction in correlation scores across all biases, demonstrating that the revealed groups are relevant to the dataset’s underlying distribution, and can be effectively utilized with specific algorithms to mitigate the model’s dependence on spurious correlations. Fig. D.1 shows the Pearson correlation scores after an epoch of training on Waterbirds, on a subset of all biases.
Figure D.1: Correlation(sample_loss, sample_to_bias similarity) under ERM/GDRO after one epoch of training on Waterbirds. Loss correlation w/ biases, ERM vs GroupDRO using groups created with the B2T partitioning method. It can be seen that, when training with ERM, loss value is highly correlated with the biases. In contrast, GroupDRO reduces the correlations, intuitively showing that biases discovered with our method are closely related to the ground truth groups of the dataset, being used as shortcuts by the model unless mitigated.

### v3

@@ Appendix D Algorithm
Output: Identified positive SCs $\mathcal{B}$.
@@
16:     $\mathcal{B}=\mathcal{B}\cup(k,b_{k})$ // Positive SCs
17:  end forReturn $\mathcal{B}$
Appendix E Loss correlation with presence of spuriously correlated concepts
In this experiment, we look at the correlations between the loss values and the concept-to-sample similarities. We compare basic ERM with GroupDRO, applied on groups, that are obtained based on our revealed SCs (and further grouped using the B2T (Kim et al., 2024) partitioning strategy).
See in Fig. 5 how for GroupDRO, the loss-to-similarity correlation significantly decrease, revealing that the model is less prone to make mistakes on the samples containing SCs. The results show a reduction in correlation scores across all SCs, demonstrating that the revealed groups are relevant to the dataset’s underlying distribution, and can be effectively utilized with specific algorithms to mitigate the model’s dependence on spurious correlations. Fig. 5 shows the Pearson correlation scores after an epoch of training on Waterbirds, on a subset of all SCs.
Figure 5: Correlation(sample_loss, sample_to_bias similarity) under ERM/GDRO after one epoch of training on Waterbirds. Loss correlation w/ biases, ERM vs GroupDRO using groups created with the B2T partitioning method. It can be seen that, when training with ERM, loss value is highly correlated with the biases. In contrast, GroupDRO reduces the correlations, intuitively showing that biases discovered with our method are closely related to the ground truth groups of the dataset, being used as shortcuts by the model unless mitigated.